OpenStudy (anonymous):

Algebra 2 Help: A group of people were given a personality test to determine if they were Type A or Type B. The results are shown in the table below: Male Female Type A 55 75 Type B 48 22 Compare P(Male | Type B) with P(Male or Type B). A) P(Male | Type B)> P(Male or Type B) B) P(Male | Type B)< P(Male or Type B) C) P(Male | Type B) = P(Male or Type B) D) There is not enough information.

OpenStudy (misty1212):

lol i will close my version

OpenStudy (misty1212):

lets compute A) P(Male | Type B)

OpenStudy (misty1212):

there are \(48+22=70\) type B of those, \(48\) are male, so P(Male | Type B) = \(\frac{48}{70}\)

OpenStudy (anonymous):

Alright.

OpenStudy (misty1212):

P(Male or Type B) is next

OpenStudy (misty1212):

there are a total of 70 type B, and another 55 males for a total of \(70+55=125\) of one or the otehr

OpenStudy (misty1212):

\[55+75+48+22=200\]so P(male or type B) is \(\frac{125}{200}\)

OpenStudy (anonymous):

I'm following so far.

OpenStudy (misty1212):

which of course reduces to \(\frac{5}{8}\)

OpenStudy (misty1212):

compare the two numbers \[\frac{24}{35}\] vs \[\frac{5}{8}\] to see which one is larger and you are done

OpenStudy (anonymous):

The 24/35 is larger. . .?

OpenStudy (misty1212):

idk i didn't check hold on

OpenStudy (misty1212):

yes, it is !

OpenStudy (anonymous):

So the answer would be A?

OpenStudy (anonymous):

P(Male | Type B)> P(Male or Type B)

OpenStudy (misty1212):

yes i believe that is the correct one

OpenStudy (anonymous):

Thank You<3